### 10 Dec 2023 v2

@@ Abstract
Deep learning has yielded remarkable outcomes in various domains. However, the challenge of requiring large-scale labeled samples still persists in deep learning. Thus, data augmentation has been introduced as a critical strategy to train deep learning models. However, data augmentation suffers from information loss and poor performance in small sample environments. To overcome these drawbacks, we propose a feature augmentation method based on shape space theory, i.e., Geodesic curve feature augmentation, called GCFA in brevity. First, we extract features from the image with the neural network model. Then, the multiple image features are projected into a pre-shape space as features. In the pre-shape space, a Geodesic curve is built to fit the features. Finally, the many generated features on the Geodesic curve are used to train the various machine learning models. The GCFA module can be seamlessly integrated with most machine learning methods. And the proposed method is simple, effective and insensitive for the small sample datasets. Several examples demonstrate that the GCFA method can greatly improve the performance of the data preprocessing model in a small sample environment.
@@ 1 Introduction
In order to solve the shortcomings of the feature augmentation methods, this paper proposes GCFA, a Geodesic curve feature augmentation method based on the shape space theory. The shape space theory [16, 17, 18], as a kind of manifold theory, mainly focuses on the geometric attributes, structure, and arrangement of the data. The theory has been widely used in image processing due to its remarkable interpretability of data and insensitivity to limited datasets [19, 20, 21, 22]. For example, Kilian et al. proposed to use shape space for sampling and projection on a highly sparse collection of three-dimensional models to generate geometrically and semantically valid samples [23]. Han et al., on the other hand, proposed an object recognition method using the shape space theory. They projected object contours into shape space and constructed Geodesic curves consistent with the diverse potential shapes of a given object type [24]. In this paper, we extract features from images by a deep learning model and then project these image features into a pre-shape space. Then, the Geodesic curve fits these features and generates new features for machine learning or deep learning training. Thus, GCFA achieves data augmentation in a small sample scenario.
@@
  • The GCFA module can generate the data flexibly and conveniently in the small sample environments.
@@
  • Empirical results demonstrate that GCFA significantly improves the performance of downstream tasks on different benchmark datasets.
Figure 1: An overview of GCFA. The image data sequentially passes through the feature extraction module, the expand module, and the project module, which converts the image data into features and projects them into the pre-shape space. In the pre-shape space, the Geodesic curves are constructed by minimizing the sum of squares of Geodesic distance. The Geodesic curves corresponding to each category are constructed. Then, new features are generated according to the categories and given to the machine learning models to achieve the effect of feature augmentation.
@@ 3 Methods
The GCFA method consists of four steps: image feature extraction, projection of the image features into the pre-shape space as features, construction of a Geodesic curve to fit the features and generation of new data along the Geodesic curve. The main steps outlined above are depicted in Figure 3.
@@ 3.1 Image feature extraction
The initial step of the GCFA method involves mapping the input data from the pixel space to the pre-shape space. Typically, the quality of features is significantly influenced by the feature extraction method. Here, we adopt the deep learning method. Complex neural network models have a tendency to extract highly discriminative features, and the utilization of high-quality features can significantly improve the performance of downstream tasks. The image features can be extracted from each layer of the neural network models. The higher layers of the network often exhibit a stronger capability to learn more intricate and comprehensive feature representations. In this paper, we employ a Vision Transformer (ViT) network [44] of which the final fully connected layer is removed as the feature extractor. The samples are fed into the pre-trained ViT network for fine-tuning, and the fine-tuning process is continued until the cross-entropy loss stabilizes. The output features from the trained ViT network represent the features in the feature space.
@@ 3.3 Construction of a Geodesic curve to fit the features
and the step is illustrated in Figure 3 (f) and Figure 3 (g). Subsequently, assign the variable $w_{i}^{*}$ to $v_{i}^{*}$ and calculate the two furthest features from $v_{i}^{*}$, as illustrated in Figure 3 (h). Repeat the steps for obtaining $w_{i}^{*}$ as previously described in Figures 3 (i) to (j) until $w_{i}^{*}$ and $v_{i}^{*}$ cease to change. Consequently, $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$ denotes the optimal Geodesic curve fitting the distribution of feature $z_{i}^{*}$, as depicted in Figure 3 (k). The detailed procedure for GCFA of features for a specific category is outlined in Algorithm  1.
@@ 3.4 Design of the loss function
A good loss function can improve the performance of the machine learning model. Generally, the cross entropy loss function is chosen since it can more sensitively reflect the differences between the predicted and true categories. In general, the number of generated features with the GCFA module is far more than the number of the original features. To balance the influence of the quantitative differences, a new loss function is proposed based on the cross entropy loss. we employ the standard cross entropy loss function, denoted as $L_{r}$, to evaluate the model’s performance on original features. For the generated features, two key control factors are introduced, i.e., the random probability factor $P_{g}$ and the influence factor $\lambda$. In each training batch, the final classifier loss function $L_{cls}$ is defined as follows:
@@
Figure 3: The process of the GCFA method. The hyper-sphere is a pre-shape space; the blue solid dots indicate the features in the pre-shape space; the solid orange dots represent features for parameter $v_{i}^{*}$; the solid green dots represent the two farthest known features $w_{i}^{0}$ and $w_{i}^{1}$ from features $v_{i}^{*}$; (a) Projection of image features to the pre-shape space; (b) Initialization of parameter $v_{i}^{*}$. From (c) to (g), the process of finding $w_{i}^{*}$; From (h) to (k), the process of finding new $v_{i}^{*}$; (l) Obtaining an optimal Geodesic curve $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$.
Algorithm 1 Geodesic Curves Feature Augmentation (GCFA)
@@ 4 Experiments
Two datasets are utilized in the subsequent experiments. First, the proposed GCFA module is combined with various advanced learning methods and the process results demonstrate the GCFA module’s effectiveness. Then, the proposed method is compared with some data augmentation methods. Finally, some ablation experiments about classifiers are executed to check the influence of the number of features generated by GCFA. The two datasets are CIFAR-10 [45] and CIFAR-100 [45]. The benchmark methods include CoMatch [46], $\Pi$-Model [47], Mean Teacher [48], Pseudo Labeling [49], MixMatch [50], VAT [51], Dash [52], CRMatch [53], UDA [54], FixMatch [55], FlexMatch [56], FreeMatch [57], AdaMatch [58], SimMatch [59], SoftMatch [60].
4.1 Efficiently performing GCFA
The GCFA module is combined with various state-of-the-art learning methods on the CIFAR-10 and CIFAR-100 datasets [45]. CIFAR-10 and CIFAR-100 are benchmark datasets comprising 50,000 training and 10,000 test images, which are in color and have a resolution of 32×32 pixels. CIFAR-100 shares the identical image set with CIFAR-10, consisting of 100 categories, with each category comprising 600 images. The training set is partitioned into labeled and unlabeled subsets, respectively. The numbers of samples per category for two experiments on the CIFAR-10 are 4 and 20, respectively, Thus, the total labeled samples are 40 and 200 for the two experiments, respectively. Similarly, for the CIFAR-100 dataset, we maintain the same sample ratios, yielding labeled samples of 400 and 2000 for the two scenarios.
@@ 4.1 Efficiently performing GCFA
In our comparative analysis, the GCFA module is integrated with different learning models of which source codes are referenced from USB [62]. For each model, we gather experimental results from the last 10 epochs which are used to calculate the average accuracy. Simultaneously, we report a 95% confidence interval. The results are shown in Table 1. Obviously, GCFA combines different models to enhance the performance of the original model. Particularly, for the 400 labeled images in CIFAR-100, GCFA achieves an average classification accuracy improvement of approximately 4%. For the labeled images in CIFAR-10, GCFA exhibits a significant improvement when utilizing 40 labeled images. However, the enhancement of GCFA is not significant if the number of training samples is large, such as 200 labeled images in CIFAR-10. This is because the backbone network is more thoroughly trained with 400 images, and its accuracy reaches 90.13%.
To describe the influence of the GCFA module combined with the neural networks, the GCFA module is combined with ViT-T, ViT-S, WRN-28-2 [63], and WRN-28-8 [63] in supervised training. As shown in Table 2, the performance of the models are significantly improved by the GCFA module. With the increase of the training samples, the influence of the GCFA module is decreased. Therefore, the GCFA module is more suitable for the small samples.
Table 1: Classification accuracy of the learning methods combined with the GCFA module on CIFAR-10 and CIFAR-100. "Supervised" denotes that only labeled data is used to train ViT-S and ViT-T, respectively. "AVG" describes the average accuracy.
@@
- +GCFA - +GCFA - +GCFA - +GCFA
@@
Table 2: Comparison of classification accuracy whether to use or not use the GCFA module.
@@
- +GCFA - +GCFA - +GCFA - +GCFA - +GCFA - +GCFA
@@
4.2 The effectiveness of GCFA in data augmentation methods
To evaluate the effectiveness of the GCFA module, we choose several representative augmentation methods which include horizontal flip, rotate, random crop, color jitter, cutout and augmix [64]. An experiment is executed by ViT-S with some data augmentation methods and GCFA for 400 labeled images in CIFAR-10. Table  3 describes the comparison of these methods whether to combine or not combine the GCFA module. From Table 3, the GCFA module can improve the classification accuracy combined with other data augmentation methods. It is noteworthy that the combination of all data augmentation methods does not achieve optimal classification accuracy, as shown in ID 8. In contrast, optimal accuracy is achieved with the combination of the methods of horizontal flip, random crop and the GCFA module.
Table 3: Comparison of the data augmentations whether to combine or not combine the GCFA module.
@@ 4.2 The effectiveness of GCFA in data augmentation methods
horizontal flip rotate random crop color jitter cutout augmix - +GCFA
@@
4.3 The influence of machine learning models on GCFA
Two experiments that choose 40 and 200 images from CIFAR-10 as training sets, respectively, are executed to understand the influence of the GCFA module on the feature classifiers. The feature extraction network model is ViT-S, and the feature classifiers for comparison include KNN [65], SVC [66], DecisionTree [67], ExtraTree [68], RandomForest [69], Bagging [70], GradientBoost [71] and MLP [72]. The results indicate that the GCFA module can improve the performance of the machine learning models, as shown in Table 4. Moreover, the MLP with the GCFA module as the highest classification accuracy. Since MLP uses the output layer weight parameters of the ViT-S.
@@ 5 Conclusion
In this paper, we propose a data augmentation method GCFA, which can provide additional information for the models of machine learning. Experimental results show that GCFA exhibits excellent performance on different datasets. In addition, the combination of the GCFA method with other data augmentation techniques can also achieve positive results. The enhanced generality and effectiveness can provide more comprehensive enhancement effect for machine learning tasks.
There are still potential improvements for the proposed method. For instance, the influence caused by isolated points. Alternatively, if image features are excessively scattered, and the Geodesic curve fails to reflect their distribution. It is more advisable to fit these points with a Geodesic surface. Improvements in these aspects will further boost the performance of GCFA and make it more applicable to various machine learning tasks.

### 25 Dec 2023 v3

@@ Abstract
Deep learning has yielded remarkable outcomes in various domains. However, the challenge of requiring large-scale labeled samples still persists in deep learning. Thus, data augmentation has been introduced as a critical strategy to train deep learning models. However, data augmentation suffers from information loss and poor performance in small sample environments. To overcome these drawbacks, we propose a feature augmentation method based on shape space theory, i.e., feature augmentation on Geodesic curve, called FAGC in brevity. First, we extract features from the image with the neural network model. Then, the multiple image features are projected into a pre-shape space as features. In the pre-shape space, a Geodesic curve is built to fit the features. Finally, the many generated features on the Geodesic curve are used to train the various machine learning models. The FAGC module can be seamlessly integrated with most machine learning methods. And the proposed method is simple, effective and insensitive for the small sample datasets. Several examples demonstrate that the FAGC method can greatly improve the performance of the data preprocessing model in a small sample environment.
@@ 1 Introduction
In order to solve the shortcomings of the feature augmentation methods, this paper proposes FAGC, a feature augmentation method on Geodesic curve based on the shape space theory. The shape space theory [16, 17, 18], as a kind of manifold theory, mainly focuses on the geometric attributes, structure, and arrangement of the data. The theory has been widely used in image processing due to its remarkable interpretability of data and insensitivity to limited datasets [19, 20, 21, 22]. For example, Kilian et al. proposed to use shape space for sampling and projection on a highly sparse collection of three-dimensional models to generate geometrically and semantically valid samples [23]. Han et al., on the other hand, proposed an object recognition method using the shape space theory. They projected object contours into shape space and constructed Geodesic curves consistent with the diverse potential shapes of a given object type [24]. In this paper, we extract features from images by a deep learning model and then project these image features into a pre-shape space. Then, the Geodesic curve fits these features and generates new features for machine learning or deep learning training. Thus, FAGC achieves data augmentation in a small sample scenario.
@@
  • The FAGC module can generate the data flexibly and conveniently in the small sample environments.
@@
  • Empirical results demonstrate that FAGC significantly improves the performance of downstream tasks on different benchmark datasets.
Figure 1: An overview of FAGC. The image data sequentially passes through the feature extraction module, the expand module, and the project module, which converts the image data into features and projects them into the pre-shape space. In the pre-shape space, the Geodesic curves are constructed by minimizing the sum of squares of Geodesic distance. The Geodesic curves corresponding to each category are constructed. Then, new features are generated according to the categories and given to the machine learning models to achieve the effect of feature augmentation.
@@ 3 Methods
The FAGC method consists of four steps: image feature extraction, projection of the image features into the pre-shape space as features, construction of a Geodesic curve to fit the features and generation of new data along the Geodesic curve. The main steps outlined above are depicted in Figure 3.
@@ 3.1 Image feature extraction
The initial step of the FAGC method involves mapping the input data from the pixel space to the pre-shape space. Typically, the quality of features is significantly influenced by the feature extraction method. Here, we adopt the deep learning method. Complex neural network models have a tendency to extract highly discriminative features, and the utilization of high-quality features can significantly improve the performance of downstream tasks. The image features can be extracted from each layer of the neural network models. The higher layers of the network often exhibit a stronger capability to learn more intricate and comprehensive feature representations. In this paper, we employ a Vision Transformer (ViT) network [44] of which the final fully connected layer is removed as the feature extractor. The samples are fed into the pre-trained ViT network for fine-tuning, and the fine-tuning process is continued until the cross-entropy loss stabilizes. The output features from the trained ViT network represent the features in the feature space.
@@ 3.3 Construction of a Geodesic curve to fit the features
and the step is illustrated in Figure 3 (f) and Figure 3 (g). Subsequently, assign the variable $w_{i}^{*}$ to $v_{i}^{*}$ and calculate the two furthest features from $v_{i}^{*}$, as illustrated in Figure 3 (h). Repeat the steps for obtaining $w_{i}^{*}$ as previously described in Figures 3 (i) to (j) until $w_{i}^{*}$ and $v_{i}^{*}$ cease to change. Consequently, $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$ denotes the optimal Geodesic curve fitting the distribution of feature $z_{i}^{*}$, as depicted in Figure 3 (k). The detailed procedure for FAGC of features for a specific category is outlined in Algorithm  1.
@@ 3.4 Design of the loss function
A good loss function can improve the performance of the machine learning model. Generally, the cross entropy loss function is chosen since it can more sensitively reflect the differences between the predicted and true categories. In general, the number of generated features with the FAGC module is far more than the number of the original features. To balance the influence of the quantitative differences, a new loss function is proposed based on the cross entropy loss. we employ the standard cross entropy loss function, denoted as $L_{r}$, to evaluate the model’s performance on original features. For the generated features, two key control factors are introduced, i.e., the random probability factor $P_{g}$ and the influence factor $\lambda$. In each training batch, the final classifier loss function $L_{cls}$ is defined as follows:
@@
Figure 3: The process of the FAGC method. The hyper-sphere is a pre-shape space; the blue solid dots indicate the features in the pre-shape space; the solid orange dots represent features for parameter $v_{i}^{*}$; the solid green dots represent the two farthest known features $w_{i}^{0}$ and $w_{i}^{1}$ from features $v_{i}^{*}$; (a) Projection of image features to the pre-shape space; (b) Initialization of parameter $v_{i}^{*}$. From (c) to (g), the process of finding $w_{i}^{*}$; From (h) to (k), the process of finding new $v_{i}^{*}$; (l) Obtaining an optimal Geodesic curve $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$.
Algorithm 1 Feature Augmentation on Geodesic Curve(FAGC)
@@ 4 Experiments
Two datasets are utilized in the subsequent experiments. First, the proposed FAGC module is combined with various advanced learning methods and the process results demonstrate the FAGC module’s effectiveness. Then, the proposed method is compared with some data augmentation methods. Finally, some ablation experiments about classifiers are executed to check the influence of the number of features generated by FAGC. The two datasets are CIFAR-10 [45] and CIFAR-100 [45]. The benchmark methods include CoMatch [46], $\Pi$-Model [47], Mean Teacher [48], Pseudo Labeling [49], MixMatch [50], VAT [51], Dash [52], CRMatch [53], UDA [54], FixMatch [55], FlexMatch [56], FreeMatch [57], AdaMatch [58], SimMatch [59], SoftMatch [60].
4.1 Efficiently performing FAGC
The FAGC module is combined with various state-of-the-art learning methods on the CIFAR-10 and CIFAR-100 datasets [45]. CIFAR-10 and CIFAR-100 are benchmark datasets comprising 50,000 training and 10,000 test images, which are in color and have a resolution of 32×32 pixels. CIFAR-100 shares the identical image set with CIFAR-10, consisting of 100 categories, with each category comprising 600 images. The training set is partitioned into labeled and unlabeled subsets, respectively. The numbers of samples per category for two experiments on the CIFAR-10 are 4 and 20, respectively, Thus, the total labeled samples are 40 and 200 for the two experiments, respectively. Similarly, for the CIFAR-100 dataset, we maintain the same sample ratios, yielding labeled samples of 400 and 2000 for the two scenarios.
@@ 4.1 Efficiently performing FAGC
In our comparative analysis, the FAGC module is integrated with different learning models of which source codes are referenced from USB [62]. For each model, we gather experimental results from the last 10 epochs which are used to calculate the average accuracy. Simultaneously, we report a 95% confidence interval. The results are shown in Table 1. Obviously, FAGC combines different models to enhance the performance of the original model. Particularly, for the 400 labeled images in CIFAR-100, FAGC achieves an average classification accuracy improvement of approximately 4%. For the labeled images in CIFAR-10, FAGC exhibits a significant improvement when utilizing 40 labeled images. However, the enhancement of FAGC is not significant if the number of training samples is large, such as 200 labeled images in CIFAR-10. This is because the backbone network is more thoroughly trained with 400 images, and its accuracy reaches 90.13%.
To describe the influence of the FAGC module combined with the neural networks, the FAGC module is combined with ViT-T, ViT-S, WRN-28-2 [63], and WRN-28-8 [63] in supervised training. As shown in Table 2, the performance of the models are significantly improved by the FAGC module. With the increase of the training samples, the influence of the FAGC module is decreased. Therefore, the FAGC module is more suitable for the small samples.
Table 1: Classification accuracy of the learning methods combined with the FAGC module on CIFAR-10 and CIFAR-100. "Supervised" denotes that only labeled data is used to train ViT-S and ViT-T, respectively. "AVG" describes the average accuracy.
@@
- +FAGC - +FAGC - +FAGC - +FAGC
@@
Table 2: Comparison of classification accuracy whether to use or not use the FAGC module.
@@
- +FAGC - +FAGC - +FAGC - +FAGC - +FAGC - +FAGC
@@
4.2 The effectiveness of FAGC in data augmentation methods
To evaluate the effectiveness of the FAGC module, we choose several representative augmentation methods which include horizontal flip, rotate, random crop, color jitter, cutout and augmix [64]. An experiment is executed by ViT-S with some data augmentation methods and FAGC for 400 labeled images in CIFAR-10. Table  3 describes the comparison of these methods whether to combine or not combine the FAGC module. From Table 3, the FAGC module can improve the classification accuracy combined with other data augmentation methods. It is noteworthy that the combination of all data augmentation methods does not achieve optimal classification accuracy, as shown in ID 8. In contrast, optimal accuracy is achieved with the combination of the methods of horizontal flip, random crop and the FAGC module.
Table 3: Comparison of the data augmentations whether to combine or not combine the FAGC module.
@@ 4.2 The effectiveness of FAGC in data augmentation methods
horizontal flip rotate random crop color jitter cutout augmix - +FAGC
@@
4.3 The influence of machine learning models on FAGC
Two experiments that choose 40 and 200 images from CIFAR-10 as training sets, respectively, are executed to understand the influence of the FAGC module on the feature classifiers. The feature extraction network model is ViT-S, and the feature classifiers for comparison include KNN [65], SVC [66], DecisionTree [67], ExtraTree [68], RandomForest [69], Bagging [70], GradientBoost [71] and MLP [72]. The results indicate that the FAGC module can improve the performance of the machine learning models, as shown in Table 4. Moreover, the MLP with the FAGC module as the highest classification accuracy. Since MLP uses the output layer weight parameters of the ViT-S.
@@ 5 Conclusion
In this paper, we propose a data augmentation method FAGC, which can provide additional information for the models of machine learning. Experimental results show that FAGC exhibits excellent performance on different datasets. In addition, the combination of the FAGC method with other data augmentation techniques can also achieve positive results. The enhanced generality and effectiveness can provide more comprehensive enhancement effect for machine learning tasks.
There are still potential improvements for the proposed method. For instance, the influence caused by isolated points. Alternatively, if image features are excessively scattered, and the Geodesic curve fails to reflect their distribution. It is more advisable to fit these points with a Geodesic surface. Improvements in these aspects will further boost the performance of FAGC and make it more applicable to various machine learning tasks.